# Error Crafting in Probabilistic Quantum Gate Synthesis

### v1

@@ Abstract
At the early stage of fault-tolerant quantum computing, it is envisioned that the gate synthesis of a general unitary gate into universal gate sets yields error whose magnitude is comparable with the noise inherent in the gates themselves. While it is known that the use of probabilistic synthesis already suppresses such coherent errors quadratically, there is no clear understanding on its remnant error, which hinders us from designing a holistic error countermeasure that is effectively combined with error suppression and mitigation. In this work, we propose that, by exploiting the fact that synthesis error can be characterized completely and efficiently, we can craft the remnant error of probabilistic synthesis such that the error profile satisfies desirable properties. We prove for the case of single-qubit unitary synthesis that, there is a guaranteed way to perform probabilistic synthesis such that we can craft the remnant error to be described by Pauli and depolarizing errors, while the conventional twirling cannot be applied in principle. Furthermore, we show a numerical evidence for the synthesis of Pauli rotations based on Clifford+T formalism that, we can craft the remnant error so that it can be eliminated up to cubic order by combining logical measurement and feedback operations. As a result, Pauli rotation gates can be implemented with T counts of $\log_{2}(1/\varepsilon)$ on average up to accuracy of $\varepsilon=10^{-9}$, which can be applied to early fault-tolerant quantum computation beyond classical tractability. Our work opens a novel avenue in quantum circuit design and architecture that orchestrates error countermeasures.
@@ Introduction
It is widely accepted that quantum computers with error correction [1, 2] can be potentially useful in various tasks such as factoring [3], quantum many-body simulation [4], quantum chemistry [5], and machine learning [6]. With the tremendous advancements in the quantum technology, the earliest experimental demonstrations of quantum error correction are realized with various codes  [7, 8, 9, 10, 11, 12, 13, 14, 15]. While such a progress towards fault-tolerant quantum computing (FTQC) is promising, quantum computers in the coming generation, the early fault-tolerant quantum computers, would still be subject to noise at the logical level, and hence the error countermeasures developed for noisy intermediate-scale quantum (NISQ) devices, in particular the quantum error mitigation (QEM) techniques [16, 17, 18, 19, 20, 21, 22, 23], will remain essential to exemplify practical quantum advantage [24, 25, 26, 27, 28, 29, 30]. Meanwhile, there is a stark difference between (early) FTQC and NISQ regimes that is characterized by the Eastin-Knill theorem [31]; the universal gate set is discrete. This indicates that the gate synthesis yields coherent error whose effect is comparable to incoherent errors triggered by noise in the hardware.
@@
In contrast to incoherent errors which inevitably necessitate entropy reduction via measurement and feedback [32, 33, 34, 35], there is no corresponding fundamental limitation for coherent errors. In fact, several previous works have found that algorithmic errors can be dealt more efficiently than naive expectation [36, 37, 38, 39, 40, 41]. It was pointed out by Refs. [36, 37] that, when one have access to a Pauli rotation exposed to over and under rotations, we can simply take probabilistic mixture of two to achieve quadratic suppression of the error. Such findings have invoked the proposal of a probabilistic gate synthesis algorithm that approximately halves the consumption of magic gates under single-qubit unitary synthesis [39, 41]. It must be noted that, these existing works merely focus on the accuracy of the probabilistic synthesis, while one must take care of the compatibility between other error countermeasures to further suppress the errors. This is prominent when one desires to perform an unbiased computation via the QEM, which requires a deep knowledge of the error channel. To our knowledge, however, there is no firm understanding on the remnant error channel nor attempt to control the profile of them.
@@
In this work, we make a first step to fill this gap by introducing a notion of error crafting, i.e., perform the probabilistic synthesis that results in a remnant error channel with a desired property, for single-qubit unitaries synthesized under the framework of Clifford+T formalism. We rigorously prove that one can design the synthesis protocol to assure the remnant error to be constrained to Pauli error, which is crucial to perform efficient error mitigation techniques [16, 17, 25]. We further provide a numerical evidence that, in the case of Pauli rotation gates, one can reliably manipulate the remnant error to be biased. This allows us to perform error detection at the logical level which achieves quadratically small sampling overhead compared to the existing lower bound of sample complexity that does not rely on mid-circuit measurement. We furthermore extend our framework to perform error crafting using CPTP channels obtained from error correction at logical layer, and show that the remnant error can be suppressed up to cubic order. As a result, we can implement Pauli rotation gates of accuracy $\varepsilon$ with T count of $\log_{2}(1/\varepsilon)$ up to $\varepsilon=10^{-9}$, which is expected to be sufficient for various medium-scale FTQC beyond state-of-the-art classical computation [27, 28, 26, 30].
@@
We remark that the operation of twirling shares the spirit of transforming a noise channel into another channel with better-known property via randomized gate compilation. However, there are two severe limitations in twirling: first, it agnostically smears out the characteristics of the channel, and second, it cannot be applied to noise maps following non-Clifford operations in general. This is in sharp contrast to our work, since our objective is to fully utilize the information on the error profile of general unitary synthesis that can be characterized completely and efficiently via classical computation.
Unitary, probabilistic, and error-crafted synthesis
@@
We aim to implement a target single-qubit unitary $\hat{U}$ with its channel representation given by $\hat{\mathcal{U}}$. We hereafter distinguish the unitaries before and after synthesis by the presence and absence of a hat. Due to the celebrated Eastin-Knill’s theorem [31], there is no fault-tolerant implementation of continuous gate sets, and therefore we cannot exactly implement the target in general. We alternatively aim to approximate $\hat{U}$ by decomposing it to a sequence of implementable operations. Such a task is called gate synthesis.
@@
One of the most common approaches for gate synthesis is the unitary synthesis, namely to decompose the target unitary into sequence of implementable unitary gate sets. For instance, unitary synthesis into the Clifford+T gate set is known to be possible with classical computation time of $O({\rm polylog}(1/\epsilon))$ for arbitrary single-qubit Pauli rotation assuming a factoring oracle [42] and $O({\rm poly}(1/\epsilon))$ for a general SU(2) gate [43, 44]. Another measure for the efficiency of synthesis is the number of T gates, or the T count, since T gates require cumbersome procedures such as magic state distillation and gate teleportation. In this regard, the existing methods saturates the information-theoretic lower bound regarding the T count of $3\log_{2}(1/\epsilon)$ except for rare exceptions that require $4\log_{2}(1/\epsilon)$ [45, 42, 43, 44].
@@
We may also consider probabilistic synthesis, which takes probabilistic mixture of several different outputs from the unitary synthesis algorithm. Concretely, we generate a set of slightly shifted unitaries $\{U_{j}\}$ which are $\epsilon$-close approximations of the target unitary $\hat{U}$ in terms of diamond distance as $d_{\diamond}(\hat{\mathcal{U}},\mathcal{U}_{j})\leq\epsilon$ (see Supplementary Note S1 for the definition of diamond distance). It is known that the probabilistic synthesis allows quadratic suppression of the error, when we appropriately generate a set of synthesized unitaries $\{\mathcal{U}_{j}\}$ and choose the probability distribution over them:
@@
(informal [41]) Let $\hat{\mathcal{U}}$ be a target unitary channel with $\{\mathcal{U}_{j}\}$ constituting its $\epsilon$-net in terms of the diamond distance. Then, one can find the optimal probabilistic compilation $\{p_{j}\}$ such that the diamond distance is bounded as
@@
As an extension of existing synthesis techniques, here we propose error-crafted synthesis, or error crafting. In short, this can be understood as constraining the remnant error of the synthesis, depending on the subsequent error countermeasure (such as QEM or error detection) one wish to perform. We formally define the remnant error channel as
@@
where $P_{a}\in\{{\rm I},{\rm X},{\rm Y},{\rm Z}\}$ and $\{\chi_{ab}\}$ gives the chi representation of the remnant error. We also allow the implementable channel to be non-unitary CPTP map as $\{\Lambda_{j}\},$ which we find to be useful when we incorporate measurement and feedback as discussed later. The aim of the crafted synthesis is to control $\{\chi_{ab}\}$ by optimizing the probability distribution $\{p_{j}\}$ for appropriately generated $\{\Lambda_{j}\}$.
Let us remark that we require the synthesis protocol to be distinct from the QEM stage, in a sense that we do not allow synthesis to increase the sampling overhead, the multiplication factor in the number of circuit execution. Also, we require that the gate complexity is nearly homogeneous even if we include feedback operations, in order to prevent a situation where parallel application of multiple synthesized channels is stalled due to a fallback scheme [46, 39].
Figure 1: Numerical analysis on error-crafted synthesis for the Pauli constraint. (a) Diamond distance $d_{\diamond}(\hat{\mathcal{U}},\mathcal{U}_{\rm synth})$ between the target single-qubit Haar random unitary $\hat{\mathcal{U}}$ and the synthesized channel $\mathcal{U}_{\rm synth}\coloneqq\sum_{j}p_{j}\mathcal{U}_{j}$. The black real line indicates the results from unitary synthesis by the Ross-Selinger algorithm, while the filled dots are results from probabilistic synthesis with shift factors of $c=2,3,5,7.$ When we increase the number of synthesized unitaries by a factor of $R=3$ for $c=7$, shown by unfilled triangles, the remnant synthesis error is suppressed below the bound in Eq. (5) and approaches the value of $\epsilon^{2}$, which matches the upper bound for the non-constrained case as in Eq. (1). The plots are averaged over 200 trials. (b) Surface plot of failure rate $p_{\rm fail}$ of probabilistic synthesis under the Pauli constraint, averaged over 200 trials for $\epsilon=10^{-4}.$ Scaling with $c$ and $R$ are shown in (c), (d), where the number of data is 10000 and 2000, respectively.
@@
While the quadratic suppression (1) by probabilistic synthesis is already beneficial for practical quantum computing, the performance of many error counteracting protocol can be enhanced when we craft the remnant error $\mathcal{E}_{\rm rem}$. For instance, let us consider the probabilistic error cancellation (PEC), which is one of the most well-known error mitigation techniques that estimates expectation values of physical observables by quasiprobabilistic implementation of error channel inversion [16, 47]. One can show that it suffices to take Clifford operations and Pauli initialization to constitute a universal set to invert arbitrary single-qubit error channels [18, 48], while the implementation is significantly simplified when $\mathcal{E}_{\rm rem}$ is a Pauli channel. This is because the PEC can be performed merely via updates on Pauli frames, which is purely done by classical processing of stabilizer measurements [25]. As another example, when we consider the rescaling technique under white-noise approximation [49, 50], which is a provably cost-optimal way of error mitigation, we can drastically improve the accuracy of the approximation when the errors are free from coherent components (see Supplementary Note S5 for details).
@@
where $\chi_{aa}\geq 0$ is required for all $a$. We emphasize that this cannot be achieved by naively applying twirling techniques, since randomization for coherent errors following non-Clifford unitary demands non-Clifford operations, which are not error-free in FTQC. Our strategy is to prepare shift unitaries $\{\mathcal{V}_{j}^{c\epsilon}\}_{j}$ that can be used to generate a set of $c\epsilon$-close unitaries such that the feasibility of the minimization problem (4) is assured even when unitary synthesis induces error of $\epsilon$. By explicitly constructing such a set of shift unitaries, we rigorously answer to the above-raised question as the following:
@@
There exist shift factor $c$, a positive number $\epsilon_{0}$, and a set of shift unitaries $\{\mathcal{V}_{j}^{c\epsilon}\}_{j=1}^{7}$ such that, for any given single-qubit unitary $\hat{\mathcal{U}}$ and for any $\epsilon\in(0,\epsilon_{0}]$, if we synthesize the shifted target unitaries to obtain $\{\mathcal{U}_{j}\}=\{\mathcal{A}(\mathcal{V}_{j}^{c\epsilon}\circ\hat{\mathcal{U}})\}$ so that $d_{\diamond}(\mathcal{V}_{j}^{c\epsilon}\circ\hat{\mathcal{U}},\mathcal{U}_{j})\leq\epsilon$ is satisfied under the synthesis algorithm $\mathcal{A}$, then there exists probabilistic mixture of $\{\mathcal{U}_{j}\}$ that induces a Pauli channel as the effective synthesis error, with the accuracy being
@@
To prove Theorem 1, we explicitly construct the shift unitaries $\{\mathcal{V}_{j}^{c\epsilon}\}_{j}$, and show that the feasibility of the solution is robust even under additional synthesis error of $\epsilon.$ Once such a solution is assured, the upper bound of Eq. (5) follows directly from the formulation via the minimization over diamond distance and the fact that $\{\mathcal{U}_{j}\}_{j=1}^{7}$ is $(c+1)\epsilon$-close to $\hat{\mathcal{U}}$ with respect to the diamond distance. Also, the lower bound follows from the fact that the diamond distance between the compiled shifted unitaries $\{\mathcal{U}_{j}\}$ and the target unitary $\hat{\mathcal{U}}$ are at least $(c-1)\epsilon$. We guide readers to Supplementary Note S6 for details of the proof.
@@
Since the theorem only states the existence of finite $c$, we rely on numerical simulation to assure the practical usage with finite $c$ values. For this purpose, we synthesize a single-qubit Haar random unitary gate by first performing the axial decomposition into three Pauli rotations, each of which passed to the synthesis algorithm proposed by Ross and Selinger [42]. As shown in Fig. 1(a), the probabilistic synthesis indeed yields error-bounded solutions that achieve quadratic suppression compared to the unitary synthesis. Meanwhile, we find that the Pauli constraint is violated when we take small shift factors $c$. This is not contradicting Theorem 1, since the theorem only states the existence of $c$ which always yield successful results. It must be noted that, as shown in Fig. 1(b) and (c), the occurrence of such a violation is suppressed exponentially by increasing the value of $c$. We observe that it suffices to take $c\sim 3.5,4.5$ to yield feasible solution with failure probability of $10^{-2},10^{-3}$, respectively, and naturally deduce that $c\sim 5.5,6.5$ is required to reach $10^{-4},10^{-5}.$ We remark that, while we have employed the axial decomposition of the SU(2) gates for the results shown here, we observe similar suppression when we utilize a direct synthesis algorithm [51].
Figure 2: Numerical analysis on error crafting for remnant errors. Here we display the distribution of error rate ratio $(q_{x},q_{y},q_{z})=(p_{x}/p,p_{y}/p,p_{z}/p)$ with constraints on error profile. The unitary synthesis is done for $\epsilon=10^{-4}$ with shift factor $c=2.0,5.0$ for depolarizing and XY constraints, respectively.
@@
Since increasing the shift factor $c$ increases the prefactor of the remnant error by $c^{2}$, we find it beneficial to introduce another practical strategy that enhances the feasibility of the solution. Here we provide a simple yet powerful technique; by generating the shift unitaries for $R$ different shift factors $c/R,2c/R,...,c$ and solving the problem with $R$-fold larger number of basis set, we can both suppress the failure of error-crafted synthesis (see Fig. 1(b) and (d)) and the remnant error. As shown in Fig. 1(a), when we take $c=7$ and $R=3$, we find that the diamond distance is suppressed as $\epsilon^{2}$ for a broad range of the synthesis accuracy $\epsilon$, which is a significant improvement from the case of $R=1$ with $\sim c^{2}\epsilon^{2}$. Note that this is extremely beneficial when we wish to eliminate such remnant error using QEM methods. To be concrete, the sampling overhead is given as $(1+4p)^{L}$ under error rate of $p$ with $L$ gates when one employs the PEC method, and hence the reduction of the remnant error from $p_{\rm rem}\sim c^{2}\epsilon^{2}$ to $p_{\rm rem}\sim\epsilon^{2}$ results in exponential difference with $L$. For instance, when we implement $\{1,2,3\}\times 10^{8}$ rotation Pauli gates with $\epsilon=10^{-5}$, the sampling overhead is reduced by factors of 6.8, 47, and 317, respectively.
@@
We remark that the number of bases with nonzero probability ($\sum_{p_{j}\neq 0}1$) is always upper-bounded by 10. Hence, the actual memory footprint of transpiled circuit is comparable to ordinary probabilistic synthesis algorithms.

### v2

# Error Crafting in Mixed Quantum Gate Synthesis

@@ Abstract
In fault-tolerant quantum computing, errors in unitary gate synthesis is comparable with noise inherent in the gates themselves. While mixed synthesis can suppress such coherent errors quadratically, there is no clear understanding on its remnant error, which hinders us from designing a holistic and practical error countermeasure. In this work, we propose that the classical characterizability of synthesis error can be exploited; remnant errors can be crafted to satisfy desirable properties. We prove that we can craft the remnant error of arbitrary single-qubit unitaries to be Pauli and depolarizing errors, while the conventional twirling cannot be applied in general. For Pauli rotation gates, in particular, the crafting enables us to suppress the remnant error up to cubic order, which results in synthesis with a T-count of $\log_{2}(1/\varepsilon)$ up to accuracy of $\varepsilon=10^{-9}$. Our work opens a novel avenue in quantum circuit design and architecture that orchestrates error countermeasures.
@@ Introduction
It is widely accepted that quantum computers with error correction Shor (1995); Gottesman (1997) can be potentially useful in various tasks such as factoring Shor (1999), quantum many-body simulation Abrams and Lloyd (1999), quantum chemistry Aspuru-Guzik et al. (2005), and machine learning Biamonte et al. (2017). With the tremendous advancements in quantum technology, the earliest experimental demonstrations of quantum error correction are realized with various codes  Sundaresan et al. (2023); Egan et al. (2021); Ryan-Anderson et al. (2021); Bluvstein et al. (2024); Livingston et al. (2022); Satzinger et al. (2021); Bluvstein et al. (2022); Zhao et al. (2022); Acharya et al. (2023). While such progress towards fault-tolerant quantum computing (FTQC) is promising, quantum computers in the coming generation, the early fault-tolerant quantum computers, would still be subject to noise at the logical level, and hence the error countermeasures developed for noisy intermediate-scale quantum (NISQ) devices, in particular the quantum error mitigation (QEM) techniques Temme et al. (2017); Li and Benjamin (2017); Endo et al. (2018); Koczor (2021a); Huggins et al. (2021); Yoshioka et al. (2022); van den Berg et al. (2022); Strikis et al. (2021), will remain essential to exemplify practical quantum advantage Piveteau et al. (2021); Suzuki et al. (2022); Yoshioka et al. (2024); Babbush et al. (2018); Lee et al. (2021); Sakamoto et al. (2023); Beverland et al. (2022). Meanwhile, there is a stark difference between (early) FTQC and NISQ regimes that is characterized by the Eastin-Knill theorem Eastin and Knill (2009); currently known FTQC schemes are based on universal gate sets that consist of discrete gates. This indicates that the gate synthesis yields coherent error whose effect is comparable to incoherent errors triggered by noise in the hardware.
@@
In contrast to incoherent errors which inevitably necessitate entropy reduction via measurement and feedback Aharonov et al. (1996); Tsubouchi et al. (2023); Takagi et al. (2023); Quek et al. (2022), there is no corresponding fundamental limitation for coherent errors. In fact, several previous works have found that algorithmic errors can be dealt with more efficiently than naive expectation Hastings (2017); Campbell (2017); Wallman and Emerson (2016); Kliuchnikov et al. (2023); Akibue et al. (2023, 2024). It was pointed out by Refs. Hastings (2017); Campbell (2017) that, when one has access to a Pauli rotation exposed to over and under rotations, we can simply take a classical mixture of two to achieve quadratic suppression of the error. Such findings have invoked the proposal of a mixed synthesis scheme that approximately halves the consumption of magic gates under single-qubit unitary synthesis Kliuchnikov et al. (2023); Akibue et al. (2024). It must be noted that, these existing works merely focus on the accuracy of the mixed synthesis, while one must take care of the compatibility between other error countermeasures to further suppress the errors. This is prominent when one desires to obtain unbiased estimator via the QEM, which requires a deep knowledge of the error channel. To our knowledge, however, there is no firm understanding on the remnant error channel nor attempt to control the profile of them.
@@
In this work, we make a first step to fill this gap by introducing a notion of error crafting, i.e., perform the mixed synthesis that results in a remnant error channel with a desired property, for single-qubit unitaries synthesized under the framework of Clifford+T formalism. We rigorously prove that one can design a synthesis protocol that ensures the remnant error is constrained to Pauli error, which is crucial to perform efficient error mitigation techniques Temme et al. (2017); Li and Benjamin (2017); Suzuki et al. (2022). We further provide numerical evidence that, in the case of Pauli rotation gates, one can reliably manipulate the remnant error to be biased. This allows us to perform error detection at the logical level which achieves quadratically small sampling overhead compared to the existing lower bound of sample complexity that does not rely on mid-circuit measurement Tsubouchi et al. (2023). We furthermore extend our framework to perform error crafting using CPTP channels obtained from error correction at the logical layer, and show that the remnant error can be suppressed up to cubic order. As a result, we can implement Pauli rotation gates of accuracy $\varepsilon$ with T-count of $\log_{2}(1/\varepsilon)$ up to $\varepsilon=10^{-9}$, which is expected to be sufficient for various medium-scale FTQC beyond state-of-the-art classical computation Babbush et al. (2018); Lee et al. (2021); Yoshioka et al. (2024); Beverland et al. (2022).
@@
We remark that the operation of twirling shares the spirit, since it also transforms a noise channel into another channel with better-known properties via randomized gate compilation. However, there are two severe limitations in twirling: first, it agnostically smears out the characteristics of the channel, and second, it cannot ensure the noise maps to be stochastic when one considers non-Clifford operations beyond the third level of Clifford hierarchy (see Table 1 and Supplementary Note S1 for details). We note that existing works on randomized compiling report suppression of the noise P. Santos et al. (2024); Odake et al. (2024); Wallman and Emerson (2016), while we are not aware of any work that guarantees the structure of the noise channel. This is in sharp contrast to our work, since our objective is to fully utilize the information on the error profile of general unitary synthesis that can be characterized completely and efficiently via classical computation.
Pauli twirl Clifford twirl Crafting
T gate $\checkmark$ Partially $\checkmark$
Z rotation Partially Partially $\checkmark$
Arbitrary $\times$ $\times$ $\checkmark$
Table 1: Applicability of noise shaping techniques for unitary synthesis error in single-qubit non-Clifford gates. The row “T gate” represents the applicability for all single-qubit gates from the third level of the Clifford hierarchy. Of course, the synthesis error is not present if one includes the T gate in the implementable operations, since there is no need for crafting. The row “Z rotation” represents the applicability for Pauli rotation gates. Here, “Partially” means that we can employ a symmetric Clifford group that commutes with the target unitary to twirl the noise within each irreducible representation but not the entire space Mitsuhashi and Yoshioka (2023); Tsubouchi et al. (2024). For instance, over- and under-rotation errors cannot be twirled; one must rely on crafting.
Results
Unitary, mixed, and error-crafted synthesis
@@ Results
We aim to implement a target single-qubit unitary $\hat{U}$ with its channel representation given by $\hat{\mathcal{U}}$. We hereafter distinguish the unitaries before and after synthesis by the presence and absence of a hat. Due to the celebrated Eastin-Knill’s theorem Eastin and Knill (2009), there is no fault-tolerant implementation of continuous gate sets, and therefore we cannot exactly implement the target in general. We alternatively aim to approximate $\hat{U}$ by decomposing it to a sequence of implementable operations. Such a task is called gate synthesis.
@@
One of the most common approaches for gate synthesis is the unitary synthesis, namely to decompose the target unitary into a sequence of implementable unitary gate sets. For instance, unitary synthesis into the Clifford+T gate set is known to be possible with classical computation time of $O({\rm polylog}(1/\epsilon))$ for arbitrary single-qubit Pauli rotation assuming a factoring oracle Ross and Selinger (2016) and $O({\rm poly}(1/\epsilon))$ for a general SU(2) gate Fowler (2011); Bocharov and Svore (2012). Another measure for the efficiency of synthesis is the number of T gates, or the T-count, since T gates require cumbersome procedures such as magic state distillation and gate teleportation. In this regard, the existing methods nearly saturates the information-theoretic lower bound regarding the T-count of $3\log_{2}(1/\epsilon)$ except for rare exceptions that require $4\log_{2}(1/\epsilon)$ Kliuchnikov et al. (2013); Ross and Selinger (2016); Fowler (2011); Bocharov and Svore (2012).
@@
We may also consider mixed synthesis, which takes a classical mixture of several different outputs from the unitary synthesis algorithm. Concretely, we generate a set of slightly shifted unitaries $\{U_{j}\}$ which are $\epsilon$-close approximations of the target unitary $\hat{U}$ in terms of diamond distance as $d_{\diamond}(\hat{\mathcal{U}},\mathcal{U}_{j})\leq\epsilon$ (see Supplementary Note S2 for the definition of diamond distance). It is known that the mixed synthesis allows quadratic suppression of the error, when we appropriately generate a set of synthesized unitaries $\{\mathcal{U}_{j}\}$ and choose the probability distribution over them:
@@
(informal Akibue et al. (2024)) Let $\hat{\mathcal{U}}$ be a target unitary channel with $\{\mathcal{U}_{j}\}$ constituting its $\epsilon$-net in terms of the diamond distance. Then, one can find the optimal weight $\{p_{j}\}$ such that the diamond distance is bounded as
@@
As an extension of existing synthesis techniques, here we propose error-crafted synthesis, or error crafting. In short, this can be understood as constraining the remnant error of the synthesis, depending on the subsequent error countermeasure (such as QEM or error detection) one wishes to perform. We formally define the remnant error channel as
@@
where $P_{a}\in\{{\rm I},{\rm X},{\rm Y},{\rm Z}\}$ and $\{\chi_{ab}\}$ gives the chi representation of the remnant error. We also allow the implementable channel to be a non-unitary CPTP map as $\{\Lambda_{j}\},$ which we find to be useful when we incorporate measurement and feedback as discussed later. The aim of the crafted synthesis is to control $\{\chi_{ab}\}$ by optimizing the probability distribution $\{p_{j}\}$ for appropriately generated $\{\Lambda_{j}\}$.
Let us remark that we require the synthesis protocol to be distinct from the QEM stage, in a sense that we do not allow synthesis to increase the sampling overhead, the multiplication factor in the number of circuit execution. Also, we require that the gate complexity is nearly homogeneous even if we include feedback operations, in order to prevent a situation where parallel application of multiple synthesized channels is stalled due to a fallback scheme Bocharov et al. (2015); Kliuchnikov et al. (2023).
Figure 1: Numerical analysis on error-crafted synthesis for the Pauli constraint. (a) Diamond distance $d_{\diamond}(\hat{\mathcal{U}},\mathcal{U}_{\rm synth})$ between the target single-qubit Haar random unitary $\hat{\mathcal{U}}$ and the synthesized channel $\mathcal{U}_{\rm synth}\coloneqq\sum_{j}p_{j}\mathcal{U}_{j}$. The black real line indicates the results from unitary synthesis by the Ross-Selinger algorithm, while the filled dots are results from mixed synthesis with shift factors of $c=2,3,5,7.$ When we increase the number of synthesized unitaries by a factor of $R=3$ (see main text for detailed definition) for $c=7$, shown by unfilled triangles, the remnant synthesis error is suppressed below the bound in Eq. (5) and approaches the value of $\epsilon^{2}$, which matches the upper bound for the non-constrained case as in Eq. (1). The plots are averaged over 200 random instances of target unitaries. (b) Surface plot of failure rate $p_{\rm fail}$ of mixed synthesis under the Pauli constraint, averaged over 200 instances for $\epsilon=10^{-4}.$ Scaling with $c$ and $R$ are shown in (c), (d), where the number of instances is 10000 and 2000, respectively.
@@
While the quadratic suppression (1) by mixed synthesis is already beneficial for practical quantum computing, the performance of many error counteracting protocols can be enhanced when we craft the remnant error $\mathcal{E}_{\rm rem}$. For instance, let us consider the probabilistic error cancellation (PEC), which is one of the most well-known error mitigation techniques that estimate expectation values of physical observables by the quasi-probabilistic implementation of error channel inversion Temme et al. (2017); van den Berg et al. (2023). One can show that it suffices to take Clifford operations and Pauli initialization to constitute a universal set to invert arbitrary single-qubit error channels Endo et al. (2018); Takagi (2021), while the implementation is significantly simplified when $\mathcal{E}_{\rm rem}$ is a Pauli channel. This is because the PEC can be performed merely via updates on Pauli frames, which is purely done by classical processing of stabilizer measurements Suzuki et al. (2022). As another example, when we consider the rescaling technique under white-noise approximation Arute et al. (2019); Dalzell et al. (2024), which is a provably cost-optimal way of error mitigation, we can drastically improve the accuracy of the approximation when the errors are free from coherent components (see Supplementary Note S6 for details).
@@
where $\chi_{aa}\geq 0$ is required for all $a$. We emphasize that this cannot be achieved by naively applying twirling techniques, since randomization for coherent errors following non-Clifford unitary demands non-Clifford operations, which are not error-free in FTQC (see Supplementary Note S1). Our strategy is to prepare shift unitaries $\{\mathcal{V}_{j}^{c\epsilon}\}_{j}$ that can be used to generate a set of $c\epsilon$-close unitaries such that the feasibility of the minimization problem (4) is guaranteed, even when unitary synthesis induces error of $\epsilon$. By explicitly constructing such a set of shift unitaries, we rigorously answer to the above-raised question as the following:
@@
There exist shift factor $c$, a positive number $\epsilon_{0}$, and a set of shift unitaries $\{\mathcal{V}_{j}^{c\epsilon}\}_{j=1}^{7}$ such that, for any given single-qubit unitary $\hat{\mathcal{U}}$ and for any $\epsilon\in(0,\epsilon_{0}]$, if we synthesize the shifted target unitaries to obtain $\{\mathcal{U}_{j}\}=\{\mathcal{A}(\mathcal{V}_{j}^{c\epsilon}\circ\hat{\mathcal{U}})\}$ so that $d_{\diamond}(\mathcal{V}_{j}^{c\epsilon}\circ\hat{\mathcal{U}},\mathcal{U}_{j})\leq\epsilon$ is satisfied under the synthesis algorithm $\mathcal{A}$, then there exists a mixture of $\{\mathcal{U}_{j}\}$ that induces a Pauli channel as the effective synthesis error, with the accuracy being
@@
To prove Theorem 1, we explicitly construct the shift unitaries $\{\mathcal{V}_{j}^{c\epsilon}\}_{j}$, and show that the feasibility of the solution is robust even under additional synthesis error of $\epsilon.$ Once such a solution is ensured, the upper bound of Eq. (5) follows directly from the formulation via the minimization over diamond distance and the fact that $\{\mathcal{U}_{j}\}_{j=1}^{7}$ is $(c+1)\epsilon$-close to $\hat{\mathcal{U}}$ with respect to the diamond distance. Also, the lower bound follows from the fact that the diamond distance between the compiled shifted unitaries $\{\mathcal{U}_{j}\}$ and the target unitary $\hat{\mathcal{U}}$ are at least $(c-1)\epsilon$. We guide readers to Supplementary Note S7 for details of the proof.
@@
Since the theorem only states the existence of finite $c$, we rely on numerical simulation to verify the practical usage with finite $c$ values. For this purpose, we synthesize a single-qubit Haar random unitary gate by first performing the axial decomposition into three Pauli rotations, each of which passed to the synthesis algorithm proposed by Ross and Selinger Ross and Selinger (2016). As shown in Fig. 1(a), the mixed synthesis indeed yields error-bounded solutions that achieve quadratic suppression compared to the unitary synthesis. Meanwhile, we find that the Pauli constraint is violated when we take small shift factors $c$. This is not contradicting Theorem 1, since the theorem only states the existence of $c$ which always yield successful results. It must be noted that, as shown in Fig. 1(b) and (c), the occurrence of such a violation is suppressed exponentially by increasing the value of $c$. We observe that it suffices to take $c\sim 3.5,4.5$ to yield feasible solution with failure probability of $10^{-2},10^{-3}$, respectively, and naturally deduce that $c\sim 5.5,6.5$ is required to reach $10^{-4},10^{-5}.$ We remark that, while we have employed the axial decomposition of the SU(2) gates for the results shown here, we observe similar suppression when we utilize a direct synthesis algorithm Morisaki et al. (2024).
@@
Since increasing the shift factor $c$ increases the prefactor of the remnant error by $c^{2}$, we find it beneficial to introduce another practical strategy that enhances the feasibility of the solution. Here we provide a simple yet powerful technique; by generating the shift unitaries for $R$ different shift factors $c/R,2c/R,...,c$ and solving the problem (4) with $R$-fold larger number of basis set, we can both suppress the failure of error-crafted synthesis (see Fig. 1(b) and (d)) and the remnant error. As shown in Fig. 1(a), when we take $c=7$ and $R=3$, we find that the diamond distance is suppressed as $\epsilon^{2}$ for a broad range of the synthesis accuracy $\epsilon$, which is a significant improvement from the case of $R=1$ with $\sim c^{2}\epsilon^{2}$. Note that this is extremely beneficial when we wish to eliminate such remnant errors using QEM methods. To be concrete, the sampling overhead is given as $(1+4p)^{L}$ under error rate of $p$ with $L$ gates when one employs the PEC method, and hence the reduction of the remnant error from $p_{\rm rem}\sim c^{2}\epsilon^{2}$ to $p_{\rm rem}\sim\epsilon^{2}$ results in exponential difference with $L$. For instance, when we implement $\{1,2,3\}\times 10^{8}$ rotation Pauli gates with $\epsilon=10^{-5}$, the sampling overhead is reduced by factors of 6.8, 47, and 317, respectively.
Figure 2: Comparison of the magnitude of remnant errors $d_{\diamond}(\mathcal{E_{\rm rem}},\mathcal{I})$. (a) Box plots are provided for the magnitude of remnant errors of 100 random instances of Z rotation gate, with unitary synthesis done for $\epsilon=10^{-4}$. The left two box plots show $d_{\diamond}(\mathcal{E}_{\rm rem},\mathcal{I})$ with and without crafting for Pauli error, and the right two box plots indicate the magnitude of the non-Pauli contribution. (b) Scaling of the remnant errors with $\epsilon.$ In both panels, shift unitaries are generated with $c=3,R=5.$
We remark that crafting is actually essential to guarantee the remnant channel to be a Pauli channel. In Fig. 2, we compare the remnant errors in crafted and uncrafted mixed synthesis. We find that, as expected, the magnitude of the remnant error is slightly smaller when we do not craft it. However, we can see there is a drastic difference in the non-Pauli component. Here, we calculate the magnitude of the error that persists even after we eliminate the Pauli component as $\min_{\mathcal{R}}d_{\diamond}(\mathcal{R}\circ\mathcal{E}_{\rm rem},\mathcal{I})$, where $\mathcal{R}$ is minimized over the set of trace-preserving maps that inverts Pauli error. We can see that the non-Pauli component in uncrafted synthesis remains as $O(\epsilon^{2})$, while in the crafted synthesis the magnitude is suppressed below $10^{-13}$. We attribute this residual to the relaxation of equality constraint into inequality, which was performed to enhance the stability of the numerical solver (see Supplementary Note S3). While we show the result for Z rotation here, we have confirmed that similar results can be obtained for general SU(2) gates as well.
@@
As another practical remark, note that the number of bases with nonzero probability ($\sum_{p_{j}\neq 0}1$) is always upper-bounded by 10. Hence, the actual memory footprint of the transpiled circuit is comparable to ordinary mixed synthesis algorithms.